sir ,i have tried to find the velocity of electron in the valence shell of hydrogen by equating the electrostatic force between electron and protons with the formula mv^2/r where m is the mass of electron and r is the atomic radius .is this method correct or not?if not ,then suggest other method to find the velocity of electron in a shell.

Grade:

1 Answers

vikas askiitian expert
13 years ago

this method is correct but u should have to use bohrs quantaization of angular momentam...

mvr = nh/2pi                         ...........1  

mv2/r = k(Z)e/r2                   ...............2

use both these equations